Who knows the game sipa? A physical education major seeks to determine whether all Filipinos are familiar with this game. Among 1,500 respondents, 358 know the game. Use 95% confidence to estimate the population proportion p and q. Interpret the sample data.


1
Expert's answer
2021-05-20T13:59:38-0400

"n=1500,p=\dfrac{358}{1500}=0.2387=23.87\%,\ q=76.13\%"


"\alpha=0.05,Z_{\frac{\alpha}{2}}=z_{0.025}=1.96"


population parameter-


"\hat{p}=p\pm Z_{0.025}\sqrt{\dfrac{pq}{n}}"


"=0.2387\pm 1.96 \sqrt{\dfrac{0.2387\times 0.7613}{1500}}\\[9pt]=0.2387\pm 0.02157\\[9pt]=(0.217,0.25027)"